OpenStudy (anonymous):

What is the value of t in the proportion? 5.6 - t / t + 3.4 = -1/4 / means over

OpenStudy (anonymous):

-1/4 * (5.6 - t) = t + 3.4 -1.4 + t/4 = t + 3.4 -3t/4 = 4.8 t = -6.4

OpenStudy (anonymous):

wait that's wrong

OpenStudy (anonymous):

-1/4 * (t+3.4) = 5.6 - t -t/4 - 0.85 = 5.7 - t 3t/4 = 6.55 t = 8.7333...